# Generalizable Prompt Tuning for Vision-Language Models

### v2

@@ 1. Introduction
In CLIP, the model is trained to associate images and corresponding texts, such that the representations of the two modalities are close in the joint embedding space (Radford et al., 2021). The pre-training process involves generating many such image-text pairs from a large corpus of images and texts. Prompt tuning involves finding the best text prompts to use for generating these pairs. This can involve experimenting with different prompts to see which ones lead to the best performance on specific tasks, such as image classification or captioning (Liu et al., 2023b). Specifically, prompt tuning can be divided into two categories:
@@
Soft prompts. Another approach is to use algorithms or other search techniques to explore the space of possible prompts and find the ones that lead to the best performance (Cho et al., 2021; Jia et al., 2022; Zhou et al., 2022b, a; Zhu et al., 2022; Yao et al., 2023). In this approach, soft prompts are often utilized, represented by the format “v1, v2, , vM, [CLASS]”, where vM denotes the learnable vector and is optimized by downstream tasks.
@@ 2.1. Vision-Language Models
Recent research has demonstrated the the necessity of multimodal learning that vision-language models (VLMs) trained on image-text pairs are more capable than those that rely solely on images. This model has received benefits from the advancements in three specific areas. Firstly, text representation learning using Transformers (Vaswani et al., 2017) provides better feature representation abilities. Secondly, large-minibatch contrastive representation learning methods (Chen et al., 2020; He et al., 2020) help the network to learn discriminative information. Finally, web-scale training datasets (Radford et al., 2021; Jia et al., 2021), which contain billions of image-text pairs for pretraining (0.4 billion pairs for CLIP (Radford et al., 2021), 1.8 billion pairs for ALIGN (Jia et al., 2021)), have been utilized. To accurately annotate image-text pairs at the billion level, unsupervised or weakly supervised learning methods (Wang et al., 2022) are employed in the raw training dataset. Specifically, ViLT (Kim et al., 2021) improves the robustness of visual and text embedding by randomly masking words in the text. Additionally, masked autoencoders (He et al., 2022) randomly mask patches of the input image to formulate a capable self-supervised learner. CLIP (Radford et al., 2021) is the representative VLM, which trains the visual encoder and visual encoder using the contrastive loss based on 400 million image-text pairs obtained from the internet, exhibiting impressive zero-shot ability. Recent studies are devoted to improve the generalization ability of VLMs, hallucination issue (Huo et al., 2024) and OOD (Zhou et al., 2022a). Our study is based on adapting pre-trained VLMs to downstream applications using CLIP, following prompt tuning methods (Zhou et al., 2022b, a; Lu et al., 2022; Yao et al., 2023; Zhu et al., 2022, 2023).
@@ 2.2. Prompt Tuning
Pre-trained VLMs have demonstrated success in various tasks, such as point cloud understanding (Zhang et al., 2022), dense prediction (Rao et al., 2022), continual learning (Thengane et al., 2022; Zheng et al., 2023), and visual understanding (Rasheed et al., 2023), among others. Prompt tuning (Liu et al., 2023b) is one of the most promising approaches to adapting pretrained VLMs to specific tasks. Originally introduced in Large Language Models (LLMs), prompt tuning involves hand-crafted instructions (or examples) for the task input to guide the LLMs to generate the appropriate output (Brown et al., 2020). Jiang et al. (Jiang et al., 2020) replaced the hand-crafted prompts with candidate prompts obtained through text mining and paraphrasing, and during the training process, they selected the optimal candidates. The gradient-based approach (Shin et al., 2020) selected the best tokens from the vocabulary that resulted in the most significant changes in gradients, based on the label likelihood. In VLMs, task-related tokens are applied to infer task-specific textual knowledge through prompt tuning. For example, CLIP (Radford et al., 2021) uses a hand-crafted template, such as “a photo of a [CLASS]”, to model the text embedding for zero-shot prediction.
@@
Our work is distinct from others in the following ways. Of the methods mentioned above, our work is most closely related to CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), ProGrad (Zhu et al., 2022), and KgCoOp (Yao et al., 2023). CoOp serves as the baseline method. Based on it, CoCoOp alleviates the overfitting issue by introducing extra conditional context vectors. Both ProGrad and KgCoOp aim to align learnable specific knowledge (i.e., learnable prompt) with general knowledge (i.e., hand-crafted prompt). ProGrad employs Kullback-Leibler (KL) divergence loss and modulates gradient updating to achieve this alignment, while KgCoOp constrains the euclidean distance between embedding distances of learnable and hand-crafted prompts. However, these two methods use hard loss functions and may not achieve optimal results for both task-specific and general abilities. In addition to the challenge of balancing task-specific and general abilities, there is still untapped potential in exploring the visual modality. Our work sets itself apart from others by meticulously addressing these crucial limitations, leading to superior outcomes in terms of the trade-off between task-specific and general abilities (as referred to Section 4).
@@ 2.3. Data Augmentation
There is a substantial amount of literature on data augmentation methods to enhance the generalization ability of neural networks. Recently, label mixing-based approaches, such as Mixup (Zhang et al., 2017), Cutmix (Yun et al., 2019), and Saliency-Mixup (Uddin et al., 2021), have been shown to significantly improve the generalization of neural networks. These mixup strategies have been successfully employed in various visual tasks, such as continual learning (Zhu et al., 2021) and long-tailed visual recognition (Yang et al., 2022). Despite their success, current vision-language prompt tuning methods have not focused on utilizing image augmentation techniques to enhance generalization ability. Interestingly, previous methods that adopted the same Mixup strategy as ours did not show significant improvements and even performed worse in some metrics, as demonstrated in Table 2. One potential reason for this outcome is that, with only learnable parameters in the text inputs, the uninformative pixels generated by Mixup (Zhang et al., 2017) could mislead existing methods and cause them to learn unexpected feature representations, even noises in the under-parameterized regime (Liu et al., 2023a). Possible solutions to this problem could involve advanced augmentation strategies (Yun et al., 2019; Uddin et al., 2021) that selectively mix informative regions of two classes, or update the visual backbone, which could potentially disturb the pre-trained parameters and cause misalignment between visual and textual information. In this work, we innovatively employ a learnable mutual information estimator to maximize the mutual information of dual views of augmented features, which enables us to balance task-specific and general abilities and effectively leverage the augmented visual modality.
@@ 2.4. Knowledge Ensemble
Knowledge distillation and sample replay are commonly used in continual learning (Zhu et al., 2021; De Lange et al., 2022) to mitigate forgetting and adapt to new knowledge. However, prompt tuning for VLMs differs from continual learning as it assumes that the VLMs have already acquired all the necessary knowledge from the pre-training phase and focus on creating a task-specific query. Furthermore, prompt tuning does not have access to the pre-trained data and does not require memory buffers to be revisited frequently. While previous methods (Zhu et al., 2022; Yao et al., 2023) for prompt tuning used hard loss function constraints to ensemble task-specific and general knowledge, these achieved sub-optimal trade-offs (as referred to Section 4). Drawing inspiration from self-supervised representation learning (Liu et al., 2023c) that enforces the network to learn invariant feature representations, we treat the embeddings of learnable and hand-crafted prompts as dual views of the text modality and encourage the network to extract shared semantic information by maximizing mutual information. Our approach softly ensembles task-specific and general information and can also accommodate augmented features.
@@ 3.2. Motivations
While CoOp-based methods are effective in adapting pre-trained CLIP to downstream tasks, they may cause biases towards the base classes with few-shot training samples, which in turn, negatively impacts generalization ability. To address this issue, Prompt-aligned Gradient (Prograd) (Zhu et al., 2022) has been proposed, which utilizes zero-shot CLIP predictions as general knowledge to guide the optimization direction. Specifically, Prograd employs Kullback-Leibler (KL) divergence between CoOp’s prediction and the zero-shot CLIP prediction:
@@ 3.3. Fusion between Textual Modal Ensemble and Visual Modal Exploration
Textual Modal Ensemble with MI Maximization. To address the first issue discussed earlier, we propose a different approach that treats the hand-crafted and learnable prompts as two complementary views of the text modality, and we maximize the Mutual Information (MI) between them to effectively extract shared semantic information. MI is a fundamental measure of the relationship between random variables (Liu et al., 2023c), and in information theory, the integration of two variables X1 and X2 can be achieved by maximizing their MI, as this ensures that the most relevant information is retained:
@@
One may ask: “What is the affecting of class-wise augmentation for CoOp-based methods and other approaches that employ hand-crafted general knowledge?” To answer this question, we perform experiments and report the results in Table 2. As KgCoOp (Yao et al., 2023) only optimizes the text embedding and does not interact with image features, we focus our analysis on CoOp (Zhou et al., 2022b) and ProGrad (Zhu et al., 2022). We find that the Mixup strategy improves the generalization of ProGrad but negatively affects the base classes to some extent. This is because Mixup mixes two samples without discrimination, leading the classifier to learn unexpected feature representations (Yun et al., 2019; Uddin et al., 2021), even noises in the under-parameterized regime (Liu et al., 2023a). ProGrad uses hand-crafted general knowledge to alleviate overfitting but fails in base classes. On the other hand, without hand-crafted general knowledge, CoOp fails in both base and new classes.
@@ 4. Experiment
Datasets. We evaluate our proposed approach against the first and third aspects, following the settings used in (Radford et al., 2021; Zhou et al., 2022b, a; Zhu et al., 2022; Yao et al., 2023), and conduct experiments on 11 datasets: ImageNet(Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) for generic object classification; OxfordPets (Parkhi et al., 2012), Flowers102(Nilsback and Zisserman, 2008), StanfordCars (Krause et al., 2013), FGVCAircraft (Maji et al., 2013) , and Food101 (Bossard et al., 2014) for fine-grained classification; UCF101 (Soomro et al., 2012) for action recognition; EuroSAT (Helber et al., 2019) for satellite image classification; SUN397 (Xiao et al., 2010) for scene classification; DTD (Cimpoi et al., 2014) for texture classification. For the second aspect, we examine domain generalization by using ImageNet and its variants, with ImageNet serving as the source domain and its variants, including ImageNetV2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Gao et al., 2022), and ImageNet-R (Hendrycks et al., 2021), as target domains. Please refer to the Appendix for further information on each dataset.
@@
Training Details. We adopt ResNet-50 (He et al., 2016) as the backbone for the image encoder, with 16 context tokens , following (Zhou et al., 2022b; Zhu et al., 2022). The same training epochs, schedule, and data augmentation settings as (Zhou et al., 2022b; Zhu et al., 2022) are used. For all experiments, we set , , and the two-layer MI estimator has hidden dimensions of 256. To ensure fairness, we compute the final performance at least three times with different seeds and average the results. All experiments are carried out on NVIDIA RTX 3090s. For more detailed training settings, please refer to the Appendix.

### v3

@@ 1. Introduction
In CLIP, the model is trained to associate images and corresponding texts, such that the representations of the two modalities are close in the joint embedding space (Radford et al., 2021). The pre-training process involves generating many such image-text pairs from a large corpus of images and texts. Prompt tuning involves finding the best text prompts to use for generating these pairs. This can involve experimenting with different prompts to see which ones lead to the best performance on specific tasks, such as image classification or captioning (Liu et al., 2023c). Specifically, prompt tuning can be divided into two categories:
@@
Soft prompts. Another approach is to use algorithms or other search techniques to explore the space of possible prompts and find the ones that lead to the best performance (Cho et al., 2021; Jia et al., 2022; Zhou et al., 2022b, a; Zhu et al., 2022a; Yao et al., 2023). In this approach, soft prompts are often utilized, represented by the format “v1, v2, , vM, [CLASS]”, where vM denotes the learnable vector and is optimized by downstream tasks.
@@ 2.1. Vision-Language Models
Recent research has demonstrated the the necessity of multimodal learning (Huo et al., 2021b, 2022b, 2022a, 2024a, 2024d; Fan et al., 2025) that vision-language models (VLMs) trained on image-text pairs are more capable than those that rely solely on images. This model has received benefits from the advancements in three specific areas. Firstly, text representation learning using Transformers (Vaswani et al., 2017) provides better feature representation abilities. Secondly, large-minibatch contrastive representation learning methods (Chen et al., 2020; He et al., 2020) help the network to learn discriminative information. Finally, web-scale training datasets (Radford et al., 2021; Jia et al., 2021), which contain billions of image-text pairs for pretraining (0.4 billion pairs for CLIP (Radford et al., 2021), 1.8 billion pairs for ALIGN (Jia et al., 2021)), have been utilized. To accurately annotate image-text pairs at the billion level, unsupervised or weakly supervised learning methods (Wang et al., 2022) are employed in the raw training dataset. Specifically, ViLT (Kim et al., 2021) improves the robustness of visual and text embedding by randomly masking words in the text. Additionally, masked autoencoders (He et al., 2022) randomly mask patches of the input image to formulate a capable self-supervised learner. CLIP (Radford et al., 2021) is the representative VLM, which trains the visual encoder and visual encoder using the contrastive loss based on 400 million image-text pairs obtained from the internet, exhibiting impressive zero-shot ability. Recent studies are devoted to improve the generalization ability of VLMs, hallucination issue (Huo et al., 2024e) and OOD (Zhou et al., 2022a). Our study is based on adapting pre-trained VLMs to downstream applications using CLIP, following prompt tuning methods (Zhou et al., 2022b, a; Lu et al., 2022; Yao et al., 2023; Zhu et al., 2022a, 2023).
@@ 2.2. Prompt Tuning
With the development of deep learning (Zhu et al., 2022b; Huo et al., 2024c; Zhu et al., 2020; Huo et al., 2020, 2021a; Li and Huo, 2024; Huo et al., 2024b; Liu et al., 2023a), pre-trained VLMs have demonstrated success in various tasks, such as point cloud understanding (Zhang et al., 2022), dense prediction (Rao et al., 2022), continual learning (Thengane et al., 2022; Zheng et al., 2023), and visual understanding (Rasheed et al., 2023), among others. Prompt tuning (Liu et al., 2023c) is one of the most promising approaches to adapting pretrained VLMs to specific tasks. Originally introduced in Large Language Models (LLMs), prompt tuning involves hand-crafted instructions (or examples) for the task input to guide the LLMs to generate the appropriate output (Brown et al., 2020). Jiang et al. (Jiang et al., 2020) replaced the hand-crafted prompts with candidate prompts obtained through text mining and paraphrasing, and during the training process, they selected the optimal candidates. The gradient-based approach (Shin et al., 2020) selected the best tokens from the vocabulary that resulted in the most significant changes in gradients, based on the label likelihood. In VLMs, task-related tokens are applied to infer task-specific textual knowledge through prompt tuning. For example, CLIP (Radford et al., 2021) uses a hand-crafted template, such as “a photo of a [CLASS]”, to model the text embedding for zero-shot prediction.
@@
Our work is distinct from others in the following ways. Of the methods mentioned above, our work is most closely related to CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), ProGrad (Zhu et al., 2022a), and KgCoOp (Yao et al., 2023). CoOp serves as the baseline method. Based on it, CoCoOp alleviates the overfitting issue by introducing extra conditional context vectors. Both ProGrad and KgCoOp aim to align learnable specific knowledge (i.e., learnable prompt) with general knowledge (i.e., hand-crafted prompt). ProGrad employs Kullback-Leibler (KL) divergence loss and modulates gradient updating to achieve this alignment, while KgCoOp constrains the euclidean distance between embedding distances of learnable and hand-crafted prompts. However, these two methods use hard loss functions and may not achieve optimal results for both task-specific and general abilities. In addition to the challenge of balancing task-specific and general abilities, there is still untapped potential in exploring the visual modality. Our work sets itself apart from others by meticulously addressing these crucial limitations, leading to superior outcomes in terms of the trade-off between task-specific and general abilities (as referred to Section 4).
@@ 2.3. Data Augmentation
There is a substantial amount of literature on data augmentation methods to enhance the generalization ability of neural networks. Recently, label mixing-based approaches, such as Mixup (Zhang et al., 2017), Cutmix (Yun et al., 2019), and Saliency-Mixup (Uddin et al., 2021), have been shown to significantly improve the generalization of neural networks. These mixup strategies have been successfully employed in various visual tasks, such as continual learning (Zhu et al., 2021) and long-tailed visual recognition (Yang et al., 2022). Despite their success, current vision-language prompt tuning methods have not focused on utilizing image augmentation techniques to enhance generalization ability. Interestingly, previous methods that adopted the same Mixup strategy as ours did not show significant improvements and even performed worse in some metrics, as demonstrated in Table 2. One potential reason for this outcome is that, with only learnable parameters in the text inputs, the uninformative pixels generated by Mixup (Zhang et al., 2017) could mislead existing methods and cause them to learn unexpected feature representations, even noises in the under-parameterized regime (Liu et al., 2023b). Possible solutions to this problem could involve advanced augmentation strategies (Yun et al., 2019; Uddin et al., 2021) that selectively mix informative regions of two classes, or update the visual backbone, which could potentially disturb the pre-trained parameters and cause misalignment between visual and textual information. In this work, we innovatively employ a learnable mutual information estimator to maximize the mutual information of dual views of augmented features, which enables us to balance task-specific and general abilities and effectively leverage the augmented visual modality.
@@ 2.4. Knowledge Ensemble
Knowledge distillation and sample replay are commonly used in continual learning (Zhu et al., 2021; De Lange et al., 2022) to mitigate forgetting and adapt to new knowledge. However, prompt tuning for VLMs differs from continual learning as it assumes that the VLMs have already acquired all the necessary knowledge from the pre-training phase and focus on creating a task-specific query. Furthermore, prompt tuning does not have access to the pre-trained data and does not require memory buffers to be revisited frequently. While previous methods (Zhu et al., 2022a; Yao et al., 2023) for prompt tuning used hard loss function constraints to ensemble task-specific and general knowledge, these achieved sub-optimal trade-offs (as referred to Section 4). Drawing inspiration from self-supervised representation learning (Liu et al., 2023d) that enforces the network to learn invariant feature representations, we treat the embeddings of learnable and hand-crafted prompts as dual views of the text modality and encourage the network to extract shared semantic information by maximizing mutual information. Our approach softly ensembles task-specific and general information and can also accommodate augmented features.
@@ 3.2. Motivations
While CoOp-based methods are effective in adapting pre-trained CLIP to downstream tasks, they may cause biases towards the base classes with few-shot training samples, which in turn, negatively impacts generalization ability. To address this issue, Prompt-aligned Gradient (Prograd) (Zhu et al., 2022a) has been proposed, which utilizes zero-shot CLIP predictions as general knowledge to guide the optimization direction. Specifically, Prograd employs Kullback-Leibler (KL) divergence between CoOp’s prediction and the zero-shot CLIP prediction:
@@ 3.3. Fusion between Textual Modal Ensemble and Visual Modal Exploration
Textual Modal Ensemble with MI Maximization. To address the first issue discussed earlier, we propose a different approach that treats the hand-crafted and learnable prompts as two complementary views of the text modality, and we maximize the Mutual Information (MI) between them to effectively extract shared semantic information. MI is a fundamental measure of the relationship between random variables (Liu et al., 2023d), and in information theory, the integration of two variables X1 and X2 can be achieved by maximizing their MI, as this ensures that the most relevant information is retained:
@@
One may ask: “What is the affecting of class-wise augmentation for CoOp-based methods and other approaches that employ hand-crafted general knowledge?” To answer this question, we perform experiments and report the results in Table 2. As KgCoOp (Yao et al., 2023) only optimizes the text embedding and does not interact with image features, we focus our analysis on CoOp (Zhou et al., 2022b) and ProGrad (Zhu et al., 2022a). We find that the Mixup strategy improves the generalization of ProGrad but negatively affects the base classes to some extent. This is because Mixup mixes two samples without discrimination, leading the classifier to learn unexpected feature representations (Yun et al., 2019; Uddin et al., 2021), even noises in the under-parameterized regime (Liu et al., 2023b). ProGrad uses hand-crafted general knowledge to alleviate overfitting but fails in base classes. On the other hand, without hand-crafted general knowledge, CoOp fails in both base and new classes.
@@ 4. Experiment
Datasets. We evaluate our proposed approach against the first and third aspects, following the settings used in (Radford et al., 2021; Zhou et al., 2022b, a; Zhu et al., 2022a; Yao et al., 2023), and conduct experiments on 11 datasets: ImageNet(Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004) for generic object classification; OxfordPets (Parkhi et al., 2012), Flowers102(Nilsback and Zisserman, 2008), StanfordCars (Krause et al., 2013), FGVCAircraft (Maji et al., 2013) , and Food101 (Bossard et al., 2014) for fine-grained classification; UCF101 (Soomro et al., 2012) for action recognition; EuroSAT (Helber et al., 2019) for satellite image classification; SUN397 (Xiao et al., 2010) for scene classification; DTD (Cimpoi et al., 2014) for texture classification. For the second aspect, we examine domain generalization by using ImageNet and its variants, with ImageNet serving as the source domain and its variants, including ImageNetV2 (Recht et al., 2019), ImageNet-Sketch (Wang et al., 2019), ImageNet-A (Gao et al., 2022), and ImageNet-R (Hendrycks et al., 2021), as target domains. Please refer to the Appendix for further information on each dataset.
@@
Training Details. We adopt ResNet-50 (He et al., 2016) as the backbone for the image encoder, with 16 context tokens , following (Zhou et al., 2022b; Zhu et al., 2022a). The same training epochs, schedule, and data augmentation settings as (Zhou et al., 2022b; Zhu et al., 2022a) are used. For all experiments, we set , , and the two-layer MI estimator has hidden dimensions of 256. To ensure fairness, we compute the final performance at least three times with different seeds and average the results. All experiments are carried out on NVIDIA RTX 3090s. For more detailed training settings, please refer to the Appendix.